Find the accumulated amount A if the principal P is invested at the interest rate of r/year for t yr. (Round your answer to the nearest cent.)
P = $44,000, r = 6.1%, t = 7.5, compounded quarterly

Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

A=44000(1+.061/4)^30*.061
A=44000(1+.01525)^1.83
A=44000(1.028)
A=$45232.00

Answer by stanbon(75887)

Find the accumulated amount A if the principal P is invested at the interest rate of r/year for t yr. (Round your answer to the nearest cent.)
P = $44,000, r = 6.1%, t = 7.5, compounded quarterly
---------------------------------------------------------
A(t) = 44000(1+(0.061/4))^(4t)
A(7.5) = 44000(1+(0.061/4)^(4*7.5)
A(7.5) = $69,285.54
========================
Cheers,
Stan H.